HETAV PATEL

Grade 11,

5) A SOLENOID HAS TWO WINDINGS, ONE ON TOP OF THE OTHER, EACH WITH THE SAME LENGTH, NO. OF TERMS AND RESISTANCE. THE WINDING ARE SUCH THERE MAGNETIC FIELD ARE IN THE THE SAME DIRECTION, COMPARE THE MAGNETIC FIELD INSIDE THE SOLENOID WHEN THE WINDING ARE CONNECTED IN a) SERIES b)PARALLEL TO THE SAME BATTERY OF NEGLIGBLE INTERNAL REISTANCE?

5) A SOLENOID HAS TWO WINDINGS, ONE ON TOP OF THE OTHER, EACH WITH THE SAME LENGTH, NO. OF TERMS AND RESISTANCE. THE WINDING ARE SUCH THERE MAGNETIC FIELD ARE IN THE THE SAME DIRECTION, COMPARE THE MAGNETIC FIELD INSIDE THE SOLENOID WHEN THE WINDING ARE CONNECTED IN a) SERIES b)PARALLEL TO THE SAME BATTERY OF NEGLIGBLE INTERNAL REISTANCE?

Grade:12th pass

1 Answers

mrinal
8 years ago
solenoid is aso known as inductor coil (L) and grouping of inductors in series and parrell are as follows 
series L =L1+L2+L3
parallel 1/L=1/L1+1/L2+1/L3